Status Report

NOAA GOES SATOPS Morning Report – May 23, 2003  Day 143 

By SpaceRef Editor
May 23, 2003
Filed under , ,


Rehearsal #5 for the S20 spacecraft successfully completed this morning. 

All DMSP operations were nominal over the past 24 hours. 


NOAA-17 rev 4724 / F at 2049Z on May 12:  G4A playback not recovered on this rev due to ground system
problems.  100% G4A recovered on rev 4725 / F. 

All POES operations were nominal over the past 24 hours. 


GOES-9 VCC verification ground system problem is being worked today.  Test commanding looks good from
current configuration and IMC load is scheduled for later today. 

At 0115Z on May 23, GOES-9 experienced a 20-degree imager cycle slip.  Engineering staff is investigating
cause and relationship to VCC issue. 

GOES-8 continues its westward drift at 1.06 deg/day.

SpaceRef staff editor.